OpenStudy (anonymous):

Does 5(-19+17) come out to -10?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5* -19= 5* 17= then add those to together so 5* -19= -95 5* 17= 85 -95 + 85 = -10 so yes your correct

OpenStudy (mayankdevnani):